a firm has total assets of $1940000 and stockholders equity is 698000. What the debt to total asset ratio?

74%

80%

64%

None of the items
Answer #1

Calculation of debt to total assets ratio:

Debt= Total assets- stockholders equity= 1940000-698000= 1242000

Debt to total assets ratio= debt/total assets= 1242000/1940000= 0.64

So correct answer is 64%